OpenStudy (anonymous):

HELP! :( The 50th term of an arithmetic sequence is 86, and the common difference is 2. Find the first three terms of the sequence

OpenStudy (anonymous):

\[t _{n}=a+(n-1)d,86=a+(50-1)*2,a=?\] this is first term next two terms are a+2,a+2+2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (refrac532):

This is an arithmetic sequence involving only whole numbers...how did you get a decimal?

OpenStudy (refrac532):

You might be thinking about a geometric sequence.

OpenStudy (refrac532):

Think about it this way: If 86 is the 50th term, and 2 is the common difference, then what would the 49th term be?

OpenStudy (refrac532):

Remember: arithmetic sequence. It only adding or subtracting to get to the next term.

OpenStudy (refrac532):

it only requires*

OpenStudy (refrac532):

Close! But you're just a little bit off.

OpenStudy (refrac532):

Try this: \[a _{n} - n(d)\] where \[a _{n}\] is the nth term, n is the term (50), and d is the common difference.

OpenStudy (anonymous):

-14!?

OpenStudy (refrac532):

Correct!

OpenStudy (refrac532):

Oh my god I am so sorry it is -12

OpenStudy (refrac532):

I miscalculated xD

OpenStudy (refrac532):

SORRY